### v2

# MedSegMamba: 3D CNN-Mamba Hybrid Architecture for Brain Segmentation

Abstract
Widely used traditional pipelines for subcortical brain segmentation are often inefficient and slow, particularly when processing large datasets. Furthermore, deep learning models face challenges due to the high resolution of MRI images and the large number of anatomical classes involved. To address these limitations, we developed a 3D patch-based hybrid CNN-Mamba model that leverages Mamba’s selective scan algorithm, thereby enhancing segmentation accuracy and efficiency for 3D inputs. This retrospective study utilized 1784 T1-weighted MRI scans from a diverse, multi-site dataset of healthy individuals. The dataset was divided into training, validation, and testing sets with a 1076/345/363 split. The scans were obtained from 1.5T and 3T MRI machines. Our model’s performance was validated against several benchmarks, including other CNN-Mamba, CNN-Transformer, and pure CNN networks, using FreeSurfer-generated ground truths. We employed the Dice Similarity Coefficient (DSC), Volume Similarity (VS), and Average Symmetric Surface Distance (ASSD) as evaluation metrics. Statistical significance was determined using the Wilcoxon signed-rank test with a threshold of P < 0.05. The proposed model achieved the highest overall performance across all metrics (DSC 0.88383; VS 0.97076; ASSD 0.33604), significantly outperforming all non-Mamba-based models (P < 0.001). While the model did not show significant improvement in DSC or VS compared to another Mamba-based model (P-values of 0.114 and 0.425), it demonstrated a significant enhancement in ASSD (P < 0.001) with approximately 20% fewer parameters. In conclusion, our proposed hybrid CNN-Mamba architecture offers an efficient and accurate approach for 3D subcortical brain segmentation, demonstrating potential advantages over existing methods. Code is available at: https://github.com/aaroncao06/MedSegMamba.
Keywords:
Biomedical Image Processing Mamba Deep Learning Semantic Segmentation Neuroimaging
1 Introduction
Subcortical brain segmentation is a significant application in medical image processing, as it enables the extraction of quantitative structural information on subcortical regions within MRI scans. This can aid in detecting and tracking morphological deficits in various neuropsychiatric conditions, including Schizophrenia [9], Major Depressive Disorder [14], and Dementia [28]. However, achieving accurate brain segmentation has remained a challenging task due to the intricate 3D structures within the brain, the large number of anatomical labels, and the substantial computational resources required to process scans at full resolution.
While manual segmentation stands as the most trusted method, it is a labor intensive and difficult task, even for experienced clinicians. Automated tools like FreeSurfer [6] have been developed to address these challenges, providing a widely accepted standard for subcortical segmentation. Despite its widespread use, FreeSurfer’s traditional methods can be slow—requiring up to hours to process a single scan—and are often sensitive to data quality issues, which limits their efficiency when dealing with large, heterogeneous datasets.
One of the leading deep learning-based alternatives to FreeSurfer is the FastSurfer pipeline [12] [13], which is capable of performing whole-brain level segmentations. For a 2.5D approach, FastSurfer aggregates three 2D fully convolutional neural networks which utilize the classic encoder-decoder structure originating from the U-Net [24]. While FastSurfer has demonstrated superior performance compared to models such as 3D U-Net [1], QuickNAT [8], and SDNet [25], its reliance on 2D models inherently limits its ability to capture the full 3D spatial dependencies of the brain’s anatomical structures.
On the other hand, 3D patch-based solutions are better suited to capture such geometries. While full 3D volume deep learning models for segmenting many classes are currently not feasible due to data and memory constraints, a patch-based approach is more computationally efficient. It also generates more training samples per subject and better captures local 3D information.
Additionally, traditional pure CNN architectures like used in FastSurfer can suffer from the local receptive fields in each layer, making them more prone to missing out on the full global context in 3D. The Vision Transformer (ViT) architecture [3] addresses this limitation by employing a self-attention mechanism—originally developed for natural language processing [27]—that achieves state-of-the-art performance in image recognition without relying on convolutions. Each self-attention layer has a global receptive field, enabling the model to extract deeper long-range spatial dependencies. For medical image segmentation tasks, hybrid CNN-Transformer architectures modeled after the U-Net have become a popular choice, showing improved generalization and performance [2] [30] [10] [23].
However, the computational cost of Transformers scales quadratically with sequence length, which leads to substantial hardware memory requirements. This makes their application to dense prediction tasks challenging, particularly for large, high-resolution biomedical images. A shifted window-based self-attention approach [17] can mitigate this issue by reducing the computational burden, enabling the use of pure Transformer models. However, it also restricts the receptive fields in each layer, which may result in the loss of global relationships that ViT was designed to capture.
Recent advancements in state space models offer a promising alternative. Mamba [7] introduces the selective scan mechanism (S6) with a hardware aware algorithm to enhance both training and inference efficiency. This allows the model to scale linearly with sequence length while maintaining state-of-the-art performance across various long-sequence processing tasks. Since its introduction, Mamba has been adapted to numerous vision tasks, consistently demonstrating competitive results compared to both Transformer and CNN-based architectures, but with improved memory and parameter efficiency. In 2D applications, Vision Mamba [34] introduces the Vim block, which incorporates a bidirectional State-Space Model (SSM) combined with positional embeddings. Meanwhile, VMamba [16] introduced the Visual State-Space (VSS) block, centered around the 2D-Selective-Scan (SS2D) module. Because the visual components in images are not sequential like language, the Vim block unravels the image in 2 ways (forward and backward) and the SS2D block unravels it along 4 traversal paths, to mitigate any unidirectional bias in the final outputs. MambaAD [11] demonstrates that processing additional traversal paths within each Mamba module can enhance performance in 2D anomaly detection tasks. However, in the realm of 3D segmentation, current methods do not fully exploit the 3D context. For example, U-Mamba [19] unravels the image in only one direction, and SegMamba’s Tri-Oriented Mamba module [32] unfolds the image along just three directions.
In this paper, we introduce MedSegMamba, a hybrid CNN-Mamba architecture that fully leverages Mamba’s selective scan algorithm for 3D inputs, aiming to enhance subcortical brain segmentation.
2 Materials and Methods
This study did not require ethical approval as it only utilized publicly available MRI scans previously acquired for studies approved by local institutional review boards, research ethics committees, or human investigation committees.
2.1 Pipeline
Our pipeline follows a 3D patch-based approach with a hybrid CNN-Mamba model. Initially, the input scans are centered and conformed to LIA orientation, followed by intensity rescaling from 0 to 1, consistent with the steps in the FastSurfer pipeline. These input volumes with dimensions 256 x 256 x 256 are cropped and padded before patch extraction. Patches are then generated with dimensions 96 x 96 x 96, with a step size of 16 voxels between consecutive patches. Each patch is sequentially fed into the model, and the output class probabilities are reconstructed to match the original input image dimensions. The predicted probabilities for each patch are combined through a voting mechanism to determine the class for each voxel, with the final values mapped to the corresponding FreeSurfer labels. This pipeline, as illustrated in Figure 1, ensures that the model can segment an entire scan into 32 classes in under 90 seconds (all 31 subcortical structures covered by FastSurferVINN excluding cortical white matter).
Figure 1: Our pipeline extracts 3D patches from the input scan, feeds them into our model, and reconstructs the output predicted label maps to generate the subcortical segmentation of the input scan.
2.2 Model Architecture
2.2.1 3D Selective Scan (SS3D)
The selected 3D scanning pattern, depicted in Figure 2, can unravel the volume along 48 unique paths. Our core SS3D module is designed to fully exploit each possible unfolded sequence, with each module processing one of 6 possible groups of 8 sequences.
First, the input volume’s axes are transposed in one of the 6 possible ways (labeled o0 to o5), determined by the assigned orientation index of the SS3D module. The eight sequences are then extracted from the volume by rotating the volume in three different ways, unfolding it along the scanning pattern, and reversing the sequences. These eight sequences are processed independently and in parallel by the S6 blocks, with their outputs reordered and merged to reconstruct the final output volume.
Each method of unraveling the volume (48 in total) is intended to train an independent S6 block to capture the unique geometries best represented by that specific sequence order.
For the S6 blocks, we increase the SSM state dimension from the standard 16 to 64. Although this adjustment slightly reduces processing speed, it enables the model to extract more detailed information about complex structures and possible patch locations, with minimal impact on parameters and memory usage. The SSM feature expansion factor is kept at 1, as the CNN encoder before the bottleneck has already expanded the feature dimension to 1024.
(a) SS3D layout
(b) Scanning pattern
Figure 2: SS3D overview
2.2.2 VSS3D block
Our 3D Visual State Space (VSS3D) block resembles a traditional Transformer block, with self-attention replaced by our SS3D module, inspired by VMamba. It consists of two residual modules. The first module includes a sequence of layer normalization, linear projection, depth-wise convolution, SiLU activation, SS3D, followed by another layer normalization and linear projection. The second residual module consists of a layer normalization followed by a multilayer perceptron (MLP). This is shown in Figure 3
Figure 3: VSS3D block
2.2.3 Overall Architecture
The overall model architecture features a 3D CNN encoder and decoder with skip connections, and a series of VSS3D blocks as the bottleneck. Passing through the encoder, four layers of residual blocks and 3 max pooling operations downsample the input patch for an encoded feature tensor. A subsequent convolution brings the channel size to 1024, before the tensor is fed into the bottleneck. The bottleneck comprises a sequence of nine VSS3D blocks, each assigned an orientation index for its inner SS3D module. The output of the bottleneck is normalized and then passed to the decoder, which reconstructs the image to its original input dimensions. Finally, a convolution operation followed by a Softmax activation function are applied to generate a 32-channel output, where each channel represents the probability of an individual class. Each residual block within the encoder and decoder layers consists of a residual connection and two sequences of 3D Convolution, Group Normalization, and Rectified Linear Unit (ReLU). This is visualized in Figure 4.
Figure 4: MedSegMamba model architecture. Only 6 VSS3D layers are shown here, but the model contains 9 in total. The last Conv1x1 and Softmax layers are also not shown.
2.3 Data
A total of 1784 T1-weighted (T1w) MRI scans were selected from a large-scale heterogeneous dataset representing a uniformly healthy population, compiled from multiple publicly available sources [5]. All scans were acquired at a resolution of 1mm x 1mm x 1mm. Selected subjects came from the Australian Imaging Biomarkers and Lifestyle Study of Ageing (AIBL) [4], Frontotemporal Lobar Degeneration Neuroimaging Initiative (NIFD), Information eXtraction from Images (IXI), Open Access Series of Imaging Studies-1 (OASIS-1) [21], Open Access Series of Imaging Studies-2 (OASIS-2) [20], Southwest University Adult life-span Dataset (SALD) [31], Southwest University Longitudinal Imaging Multimodal Brain Data Repository (SLIM) [15], Parkinson’s Progression Markers Initiative (PPMI) [22], SchizConnect (SchizConnect) [29], and Consortium for Reliability and Reproducibility (CoRR) [35].
Figure 5: Age and Gender Distributions for each Dataset
The dataset was partitioned into training, validation, and test sets with a roughly 3:1:1 ratio. Specifically, the training set included 1076 scans, the validation set contained 345 scans, and the test set comprised 363 scans. A balanced age and gender distribution was achieved across the datasets, as shown in Figure 5. Ground truth segmentations were generated using FreeSurfer, and the input T1w scans were preprocessed with skull-stripping and intensity normalization.
2.4 Model Training
The described model was trained and evaluated on a 24 GB NVIDIA Quadro 6000 GPU. We employed the AdamW optimizer in conjunction with a Cosine Annealing Warm Restarts learning rate scheduler [18]. The loss function used was a combination of Dice Loss and Weighted Cross Entropy during the first epoch, followed by Dice Loss alone in subsequent epochs. Training was conducted on 27 patches per scan (three steps per dimension), utilizing gradient accumulation to simulate each image as a single batch, accounting for the variability in class presence across patches. A dropout rate of 0.1 and a drop path rate of 0.3 were applied within the bottleneck.
2.5 Model Evaluation and Statistical Test
To evaluate the efficacy of our novel SS3D module over other 3D mamba-based vision modules for latent space learning, we also trained a model identical to our MedSegMamba architecture, except substituting the first residual module in each VSS3D block with SegMamba’s Tri-oriented Mamba module. The resulting block is identical to SegMamba’s Tri-oriented Spatial Mamba block except excluding the Gated Spatial Convolution module, in order to focus specifically on evaluating the core Mamba modules. This model, referred to as SegMambaBot, was trained following the same procedure as MedSegMamba.
MedSegMamba was also evaluated against a retrained TABSurfer and pretrained FastSurferVINN (obtained from the FastSurfer GitHub repository) using 363 FreeSurfer segmentations as ground truths. Both SegMambaBot and TABSurfer were trained using 3D patch-based approaches with input patch sizes identical to those used in MedSegMamba.
The evaluation metrics included the Dice Similarity Coefficient (DSC), Volume Similarity (VS) [26], and Average Symmetric Surface Distance (ASSD) [33], assessing both the overall similarity of the segmentations and the contour quality relative to the ground truth. The Wilcoxon signed-rank test was employed to assess the significance of improvements in MedSegMamba’s performance over the other models for each metric. Statistical significance was set to P < 0.05.
3 Results
Figure 6: 3D renderings for a sample segmented by each method.
The average metrics from evaluating MedSegMamba, SegMambaBot, TABSurfer, and FastSurferVINN against the FreeSurfer-generated ground truths are displayed in Table 1. MedSegMamba consistently achieved high scores in Dice Similarity Coefficient (DSC), Volume Similarity (VS), and Average Symmetric Surface Distance (ASSD) across all datasets, demonstrating superior overall performance. MedSegMamba’s improvement was statistically significant across all metrics, except for the DSC and VS metrics when compared to SegMamba (P-values of 0.114 and 0.425, respectively).
All three 3D patch-based methods also significantly outperformed the pretrained 2.5D FastSurfer benchmark. Furthermore, both Mamba-based models significantly outperformed TABSurfer, their transformer-based counterpart.
Dataset Model DSC ↑ VS ↑ ASSD ↓
AIBL MedSegMamba 0.89593±0.009 0.97406±0.006 0.29069±0.042
SegMambaBot 0.89580±0.009 0.97396±0.006 0.29038±0.040
TABSurfer 0.89094±0.010 0.97082±0.007 0.30257±0.044
FastSurfer 0.87856±0.015 0.96485±0.009 0.33453±0.059
CoRR MedSegMamba 0.88635±0.020 0.97123±0.008 0.31922±0.078
SegMambaBot 0.88627±0.020 0.97040±0.008 0.32062±0.077
TABSurfer 0.87949±0.020 0.96765±0.009 0.33669±0.074
FastSurfer 0.86607±0.027 0.95989±0.016 0.38024±0.104
IXI MedSegMamba 0.86773±0.023 0.96485±0.010 0.41766±0.102
SegMambaBot 0.86729±0.023 0.96460±0.010 0.42425±0.106
TABSurfer 0.85877±0.027 0.96023±0.011 0.43849±0.108
FastSurfer 0.81217±0.034 0.93340±0.018 0.61388±0.140
NIFD MedSegMamba 0.90060±0.007 0.97734±0.006 0.26797±0.030
SegMambaBot 0.90026±0.008 0.97745±0.005 0.27077±0.032
TABSurfer 0.89260±0.008 0.97361±0.005 0.29188±0.035
FastSurfer 0.88791±0.008 0.97240±0.005 0.30494±0.030
OAS1 MedSegMamba 0.88954±0.012 0.97351±0.006 0.30856±0.050
SegMambaBot 0.88977±0.012 0.97387±0.007 0.30839±0.050
TABSurfer 0.88305±0.011 0.96808±0.005 0.32129±0.041
FastSurfer 0.87504±0.010 0.96240±0.005 0.34141±0.047
OAS2 MedSegMamba 0.88680±0.013 0.97017±0.009 0.31079±0.047
SegMambaBot 0.88670±0.012 0.97048±0.008 0.31382±0.047
TABSurfer 0.88247±0.013 0.96611±0.009 0.32258±0.047
FastSurfer 0.87984±0.013 0.96454±0.008 0.32390±0.049
PPMI MedSegMamba 0.89373±0.007 0.97345±0.005 0.29935±0.037
SegMambaBot 0.89333±0.007 0.97422±0.005 0.30096±0.036
TABSurfer 0.88941±0.008 0.97158±0.006 0.30483±0.032
FastSurfer 0.87892±0.008 0.96673±0.006 0.32847±0.033
SALD MedSegMamba 0.87785±0.028 0.96998±0.009 0.35457±0.094
SegMambaBot 0.87795±0.028 0.97015±0.009 0.35236±0.093
TABSurfer 0.87046±0.027 0.96560±0.009 0.37473±0.092
FastSurfer 0.84189±0.021 0.94426±0.014 0.48194±0.089
Schiz MedSegMamba 0.88204±0.011 0.97118±0.007 0.33188±0.044
SegMambaBot 0.88205±0.011 0.97082±0.008 0.33257±0.042
TABSurfer 0.87613±0.011 0.96747±0.007 0.34891±0.053
FastSurfer 0.83832±0.022 0.94248±0.015 0.48469±0.094
SLIM MedSegMamba 0.88692±0.006 0.97101±0.006 0.30611±0.024
SegMambaBot 0.88740±0.007 0.97169±0.006 0.30679±0.025
TABSurfer 0.88254±0.006 0.96972±0.005 0.31617±0.022
FastSurfer 0.85483±0.012 0.94886±0.008 0.42471±0.051
Overall MedSegMamba 0.88383±0.021 0.97076±0.009 0.33604±0.086
SegMambaBot 0.88372±0.021 0.97068±0.009 0.33761±0.087
TABSurfer 0.87711±0.022 0.96684±0.009 0.35264±0.088
FastSurfer 0.85350±0.035 0.95229±0.018 0.43554±0.143
Table 1: Comparing MedSegMamba, SegMambaBot, TABSurfer, and FastsurferVINN metrics across datasets. Bold text indicates superior performance. ↑ indicates that higher numbers correspond to better performance and ↓ indicates that lower numbers correspond to better performance.
Views of 3D renderings and 2D slices of segmentations produced by each method on a sample subject are shown in Figure 6 and Figure 7. Although FreeSurfer’s atlas-based method yielded the noisiest segmentation, all deep learning methods produced smoother contours. FastSurfer’s 2.5D method retained the most noise from the FreeSurfer output, whereas the 3D patch-based methods achieved similarly smooth segmentations, differing only in minor areas for this sample. However, SegMamba exhibited slight under-segmentation in certain regions, and TABSurfer demonstrated less consistency across different scans.
While both Mamba-based models showed competitive performance, MedSegMamba demonstrated slight improvements over SegMambaBot, and has approximately 20% fewer parameters while requiring similar GPU memory during inference (1.988 GiB for MedSegMamba vs. 1.932 GiB for SegMambaBot).
As shown in Figure 8, MedSegMamba demonstrates superior performance while remaining parameter efficient compared to the other 3D patch-based methods. While FastSurfer’s fully convolutional architecture requires very few parameters, each of its three sub-models requires more than double the GPU memory during inference (just under 5 GiB) compared to the 3D models, all of which operate with under 2 GiB of GPU memory.
Figure 7: 2D slices of a sample segmented by each method.
Figure 8: Plot of millions of parameters versus overall dice score for each model.
4 Discussion
This study introduces the SS3D module and VSS3D block, an extension of Mamba’s selective scan operation tailored for complex 3D image processing tasks. Integrated within the MedSegMamba hybrid CNN-Mamba architecture, the VSS3D block demonstrates robust latent-space learning capabilities, outperforming existing traditional and deep learning approaches across multiple datasets.
The transition from a Transformer bottleneck to a Mamba-based bottleneck enhances memory efficiency, enabling the use of larger convolutional layers on identical hardware. This enhancement boosts local feature extraction capabilities while preserving global context extraction, resulting in both SegMambaBot and MedSegMamba outperforming TABSurfer. Comparing SegMambaBot and MedSegMamba, the SS3D module allows MedSegMamba to attain comparable overall performance in terms of DSC and VS metrics but offers superior region boundary delineation as reflected by the ASSD, all while utilizing significantly fewer parameters than the Tri-oriented Mamba module. Increasing the number of selective scan operations to process variously unraveled sequences can achieve performance parity with larger modules that process fewer sequences.
Compared to FastSurfer’s standard 2.5D approach, our 3D inputs preserve more intricate spatial relationships within the continuity of anatomy compared to 2D slices. The 3D strided reconstruction process, which employs voting on the predicted class, further mitigates noise artifacts present in the ground truth. This improvement is observed qualitatively when comparing the noisier segmentations of FastSurfer to the smoother outputs of the 3D patch-based models.
4.1 Limitations
The primary limitation of this approach is that MedSegMamba exhibits slower performance than SegMambaBot, despite its lower parameter count. However, this issue can be mitigated by increasing the step size during the strided reconstruction process from 16 to 32, resulting in a roughly 4x speedup, reducing the processing time per scan to around 22 seconds with minimal impact on performance. When both using a step size of 32, MedSegMamba is only a few seconds slower than SegMambaBot per scan, making them roughly comparable in terms of overall efficiency while retaining the superior performance afforded by the SS3D module.
4.2 Conclusion
These results showcase the advantages of our hybrid architecture combined with a 3D patch-based approach. Future research should explore the application of our SS3D module across different architectures, including its integration at various latent space levels throughout the encoder, as seen in SegMamba’s original design. Ablation studies focused on determining the optimal number of scanning directions or unraveling methods for 3D segmentation may further enhance our model’s performance.
5 Acknowledgments
No funding was received for conducting this study and there are no relevant financial or non-financial interests to disclose.